How much water must be evaporated from 5 liters of a 30% salt solution to obtain a 50% salt solution?
Table and show work

Found 2 solutions by richard1234, robertb:
Answer by richard1234(7193)

We have 5 liters of a 30% salt solution: this means that there is 1.5 liters salt and 3.5 liters water.
To make the solution 50% salt, the amount of salt needs to be equal to the amount of water. Thus, there should be 1.5 liters of water remaining, and two liters of water should be evaporated.

Answer by robertb(5830)

0.30*5 - 0.00x = 0.50(5-x),
1.5 = 2.5 - 0.50x,
-1 = -0.50x,
x = 2 liters, the amount of water that should be evaporated.